The temperature of an ideal gas remains constant while the absolute pressure changes from 103.4 kPaa to 827.2 kPaa. (a) If the initial volume is 80 liter, what is the final volume? (b) for 160 gram of the gas, determine the change of density.


1
Expert's answer
2022-05-06T15:34:22-0400

a)

"V_2=V_1\frac{p_1}{p_2}=10~l,"

b)

"∆\rho=\frac m{V_2}-\frac m{V_1}=14~\frac gl."